# references.bib

### v2

@@ Abstract
This paper proposes a novel control method for an autonomous wheel loader, enabling time-efficient navigation to an arbitrary goal pose. Unlike prior works that combine high-level trajectory planners with Model Predictive Control (MPC), we directly enhance the planning capabilities of MPC by integrating a cost function derived from Actor-Critic Reinforcement Learning (RL). Specifically, we train an RL agent to solve the pose reaching task in simulation, then incorporate the trained neural network critic as both the stage and terminal cost of an MPC. We show through comprehensive simulations that the resulting MPC inherits the time-efficient behavior of the RL agent, generating trajectories that compare favorably against those found using trajectory optimization. We also deploy our method on a real wheel loader, where we successfully navigate to various goal poses.
@@ I Introduction
We propose an alternative control approach which omits the use of a high-level planner, and improves the planning capabilities of the MPC directly. Similar to [reiter2024ac4mpc], we leverage the fact that a trained Reinforcement Learning (RL) critic encodes all of the information needed to plan approximately optimal behavior. Thus, by integrating a trained RL critic as an MPC cost function, we enable long-term planning even with a prediction horizon suitable for real-time control.
@@
  • We demonstrate the effectiveness of our solution on a real wheel loader, where the trained RL actor is unsafe for direct use. A comprehensive study, supported by simulations, shows that our method reaches goal poses faster than a baseline trajectory optimization routine.
Compared to prior work on RL wheel loader control [10260481, sardarmehni2023path], we support a broader range of practical applications by enabling navigation to an arbitrary goal pose, instead of only executing specific maneuvers. Our approach also enhances safety by enforcing constraints through MPC. Unlike the Actor-Critic MPC in [reiter2024ac4mpc], we use a Lyapunov-based RL algorithm [wang2023policyoptimizationmethodoptimaltime] to train a critic that serves as a sampling-based Lyapunov function, providing a mean cost stability guarantee. Moreover, we use the critic not only as a terminal cost in our MPC, but also derive a suitable stage cost from it. Lastly, we propose the use of a gradient penalty during RL training to enhance the MPC optimization landscape.
@@ IV Methodology
where both the states $x$ and controls $u$ are subject to constraints. Our goal is to drive the system to a goal state $x_{g}$ and maintain it there indefinitely. To achieve this, we use the description of the nonlinear system to construct a goal-conditioned MDP, where we assign $s=x$, $a=u$, $g=x_{g}$, and define the state transitions of the MDP as deterministic using $s_{t+1}=f(s_{t},a_{t})$. We then approximate the optimal solution of the MDP through a Lyapunov-based Actor-Critic RL algorithm [wang2023policyoptimizationmethodoptimaltime], which trains a policy that stabilizes the MDP system according to Definition III.1. However, while the RL policy is able to solve the control task in simulation, it is unable to satisfy actuator limits and other constraints, making it unsuitable for real world deployment.
@@ IV-A Wheel loader kinematic model
To ensure the kinematic model remains valid, we rely on the following state constraints:
@@ IV-B Reinforcement Learning environment design
Importantly, the use of $p=0.25$ in (9) makes the cost behave in a sparse manner, i.e. giving a nearly constant cost to states which have not converged to the goal. This ensures that the optimal solution of the associated MDP is to reach the goal as quickly as possible, since that is the only way to reduce the magnitude of the cost in a significant way. Other norms, such as the Euclidean norm may encourage solutions to initially approach the goal as fast possible, but since the cost values begin to diminish very rapidly when getting closer, there is not enough consistent pressure for time-optimal convergence.
@@ IV-C Actor and critic neural networks
where $Q_{\psi}(s,a,g)$ is a feedforward neural network parameterized by $\psi$. This ensures positive outputs, which is necessary to satisfy (3a) due to the choice of $c(s,a,g)$. We also include an input encoder in both networks, which converts all positions into relative positions and encodes all heading angles using their sine and cosine values.
@@ IV-E Model Predictive Control problem formulation
Our MPC, to be described below, is formulated as the following multiple-shooting optimization problem:
@@ V-A Experimental setup
We train our RL agent using PyTorch [gardner2021gpytorchblackboxmatrixmatrixgaussian], and base our implementation on stable-baslines3 [stable-baselines3]. To help learn from the sparse cost, we use Hindsight Experience Replay (HER) [andrychowicz2018hindsightexperiencereplay] as our replay buffer $\mathcal{D}$. The actor and critic networks consist of layers with $(48,96,144,96,48)$ hidden units, employing the $\mathrm{SoftPlus}$ activation function. The RL agent is trained until convergence to $\lambda_{l}=0.8$, indicating that the critic is a sampling-based-Lyapunov function.
We formulate the MPC problem in (20) using CasADi [Andersson2019] and Acados [acados], using L4CasADi [salzmann2023learningcasadidatadrivenmodels, Salzmann_2023] to integrate the neural network critic. The resulting optimization problem is solved using a SQP-RTI scheme, where we utilize the HPIPM [frison2020hpipmhighperformancequadraticprogramming] QP solver. The solver is warm started using the solution from the previous iteration. During our our experiments we use $N=10$ and $\Delta t=200\ ms$.
@@ V-B Baseline trajectory optimization
Figure 5: The three highlighted scenarios. (a) Short loading cycle (b) Compact 180-degree turn (c) Navigation through multiple obstacles. Scenarios (a) and (b) were evaluated in the real world, while scenario (c) was conducted using simulations. The time $t$ signifies the first time instant when $||x-g||<0.1$. For scenario (c) we showcase trajectories for both $N=10$ and $N=20$ to highlight the dependence on a sufficiently long prediction horizon.
@@ V-C Highlighted scenarios
As shown in Figure 5, the Actor-Critic MPC closely follows the convergence time of the baseline in scenario (a) and slightly outperforms it in scenario (b). The velocity tracking error for scenarios (a) and (b) is illustrated in Figure 4, highlighting that the Actor-Critic MPC achieves this competitive performance despite the presence of actuator delays and velocity tracking errors, which do not impact the ideal baseline trajectories computed using (27). The difference in execution time is also significant: the baseline trajectory optimization takes over 5 seconds on a desktop AMD Ryzen 3900x CPU, while each iteration of the Actor-Critic MPC problem is solved in less than 100 milliseconds on the much weaker CPU of the Jetson. However, using $N=20$ for scenario (c) increases the MPC solve time to 200-300 milliseconds, which prevents real-world use.
V-D Supplementary analysis
@@
Improvement over baseline 23.76 % 15.83 %
@@ VI Conclusion and future work
We present a novel control scheme based on the Actor-Critic MPC framework, and apply it to solve a wheel loader pose reaching task in a time-efficient manner. Simulated experiments show that the trajectories generated by our method compare favourably against those found using nonlinear trajectory optimization. Our approach also demonstrates robustness by succeeding in real-world tests despite the presence of actuator delay and velocity tracking errors.
@@
While our empirical experiments indicate that using a Lyapunov-based RL algorithm along with a gradient penalty is advantageous, we intend to investigate their significance more rigorously in future work. Additionally, exploring Control Barrier Functions (CBFs) to shorten the required prediction horizon for obstacle avoidance, similar to [zeng2021safetycriticalmodelpredictivecontrol], is another promising direction for future research.

### v3

@@ Abstract
This paper proposes a novel control method for an autonomous wheel loader, enabling time-efficient navigation to an arbitrary goal pose. Unlike prior works which combine high-level trajectory planners with Model Predictive Control (MPC), we directly enhance the planning capabilities of MPC by incorporating a cost function derived from Actor-Critic Reinforcement Learning (RL). Specifically, we first train an RL agent to solve the pose reaching task in simulation, then transfer the learned planning knowledge to an MPC by incorporating the trained neural network critic as both the stage and terminal cost. We show through comprehensive simulations that the resulting MPC inherits the time-efficient behavior of the RL agent, generating trajectories that compare favorably against those found using trajectory optimization. We also deploy our method on a real-world wheel loader, where we demonstrate successful navigation in various scenarios.
@@ I Introduction
We propose an alternative control approach which omits the use of a high-level planner, and improves the planning capabilities of the MPC directly. Similar to [reiter2024ac4mpc], we leverage the fact that a Reinforcement Learning (RL) critic learns to encode all of the information needed to plan approximately optimal behavior. Thus, by integrating a trained RL critic as an MPC cost function, we enable long-term planning even with a prediction horizon suitable for real-time control.
@@
  • We demonstrate the effectiveness of our solution on a real wheel loader, where the RL actor would be unsafe for direct use. A comprehensive study, supported by simulations, shows that our method reaches goal poses faster than a baseline trajectory optimization routine.
Compared to prior work on RL wheel loader control [10260481, sardarmehni2023path], we support a broader range of applications by enabling navigation to an arbitrary goal pose, instead of only executing specific maneuvers. Our approach also enhances safety by enforcing constraints through MPC. Unlike the conventional Actor-Critic MPC in [reiter2024ac4mpc], we use a Lyapunov-based RL algorithm [wang2023policyoptimizationmethodoptimaltime] to train a critic that serves as a sampling-based Lyapunov function, providing a mean cost stability guarantee. Moreover, we use the critic not only as a terminal cost in our MPC, but also derive a suitable stage cost from it. Lastly, we propose the use of a gradient penalty during RL training to enhance the MPC optimization landscape.
@@ IV Methodology
where both the states $x$ and controls $u$ are subject to constraints. Our goal is to drive the system to a goal state $x_{g}$ and maintain it there indefinitely. To achieve this, we use the description of the nonlinear system to construct a goal-conditioned MDP, where we assign $s=x$, $a=u$, $g=x_{g}$, and define the state transitions of the MDP as deterministic using $s_{t+1}=f(s_{t},a_{t})$. We then approximate the optimal solution of the MDP through a Lyapunov-based Actor-Critic RL algorithm [wang2023policyoptimizationmethodoptimaltime], which trains a policy that stabilizes the MDP system according to Definition III.1. However, while the RL policy is able to solve the control task in simulation, it is unable to take into account actuator limits and other constraints, which complicates real world deployment.
@@ IV-A Wheel loader kinematic model
To ensure the simplified kinematic model remains valid, we enforce the following state constraints:
@@ IV-B Reinforcement Learning environment design
Importantly, the use of $p=0.25$ in (9) makes the cost behave in a sparse manner, i.e. giving a nearly constant cost to states which have not converged to the goal. This ensures that the optimal solution of the associated MDP reaches the goal as quickly as possible, since that is the only way to significantly reduce the magnitude of the cost. Other norms, such as the squared Euclidean norm may encourage solutions to initially approach the goal as fast possible, but since the cost values diminish very rapidly when getting closer, there is not enough consistent incentive for time-optimal convergence.
@@ IV-C Actor and critic neural networks
where $Q_{\psi}(s,a,g)$ is a feedforward neural network parameterized by $\psi$. This ensures positive outputs, which is necessary to satisfy (3a) due to the choice of $c(s,a,g)$. We also include an encoder in both networks, which converts all positions into relative positions with respect to the goal and encodes heading angles using sine and cosine values.
@@ IV-E Model Predictive Control problem formulation
Our nonlinear MPC, to be described below, is formulated as the following multiple-shooting optimization problem:
@@ V-A Experimental setup
We train our RL agent using PyTorch [gardner2021gpytorchblackboxmatrixmatrixgaussian], and base our implementation on stable-baslines3 [stable-baselines3]. The actor and critic networks consist of layers with $(48,96,144,96,48)$ hidden units, employing the $\mathrm{SoftPlus}$ activation function. The RL agent is trained until convergence to $\lambda_{l}=0.8$, indicating that the critic is a sampling-based-Lyapunov function.
We formulate the MPC problem in (20) using CasADi [Andersson2019] and Acados [acados], using L4CasADi [salzmann2023learningcasadidatadrivenmodels, Salzmann_2023] to integrate the neural network critic. The resulting optimization problem is solved using a SQP-RTI scheme, where we utilize the HPIPM [frison2020hpipmhighperformancequadraticprogramming] QP solver. The solver is warm started using the solution from the previous iteration. During our experiments we use $N=10$ and $\Delta t=0.2\ s$, which were chosen to balance the trade-off between real-time capability, length of the prediction horizon and discretization error.
@@ V-B Baseline trajectory optimization
Figure 5: The three highlighted scenarios. (a) Short loading cycle (b) Compact 180-degree turn (c) Navigation through multiple obstacles. Scenarios (a) and (b) were evaluated in the real world, while scenario (c) was conducted using simulations. The time $t$ signifies the first time instant when $||x-g||<0.1$. For scenario (c) we illustrate trajectories for both $N=10$ and $N=20$ to highlight the dependence on a sufficiently long prediction horizon.
@@ V-C Highlighted scenarios
As shown in Figure 5, the Actor-Critic MPC closely matches the baseline’s convergence time in scenario (a) and slightly outperforms it in the more challenging scenario (b). The velocity tracking error for scenarios (a) and (b) is illustrated in Figure 4, which highlights that the Actor-Critic MPC achieves this competitive performance despite the presence of minor actuator delays and velocity tracking errors. The difference in execution time is also significant: the baseline trajectory optimization takes over 5 seconds on a desktop AMD Ryzen 3900x CPU, while each iteration of the Actor-Critic MPC problem is solved in less than 100 milliseconds on the much weaker CPU of the Jetson. However, using $N=20$ for scenario (c) increases the MPC solve time to 200-300 milliseconds, which prevents successful real-world use in obstacle-rich environments.
V-D Simulation study
@@
Improvement over baseline 23.80 % 23.74 %
@@ VI Conclusion and future work
We present a novel control scheme based on the Actor-Critic MPC framework, and apply it to solve a wheel loader pose reaching task in a time-efficient manner. Our simulated experiments show that the trajectories generated by our method compare favorably against those found using nonlinear trajectory optimization. Our approach also demonstrates some robustness by succeeding in real-world tests despite the presence of actuator delay and velocity tracking errors.
@@
Although our empirical experiments indicate that using a Lyapunov-based RL algorithm along with a gradient penalty is advantageous, their significance needs to be studied more rigorously in future work. Additionally, exploring Control Barrier Functions (CBFs) to shorten the required prediction horizon for obstacle avoidance, similar to [zeng2021safetycriticalmodelpredictivecontrol, abdi2024model], is another promising direction for future research.